OpenStudy (anonymous):

HELP!!! 2x + 1 = RADICAL (3x^2 + 5x = 3)

OpenStudy (anonymous):

spamming not cool

OpenStudy (anonymous):

not spamming - need help!

OpenStudy (anonymous):

\[2x+1=\sqrt{3x^2+5x}=3\] This is what your problem looks like to me , but I do not think this is right

OpenStudy (anonymous):

not quite the last term should be +3 under the radical sign - thanks for the clarification!

OpenStudy (anonymous):

I got x= -1 and x = 2, but which one works? and am I close at all?

OpenStudy (anonymous):

Well to see which one works you would just plug those back in.

OpenStudy (anonymous):

so I get x = 5 correct?

OpenStudy (anonymous):

Yes that is correct because it works for both sides.

OpenStudy (anonymous):

x=2 worked for me!

OpenStudy (anonymous):

The x=2 is what I meant worked, so yes you did great. Looks like you knew how to do this all along. :)